Better weather today as we clear topsoil from locus E500 - warmer, but more haze. The probably LB terrace wall which was highly visible from the surface continues to be exposed on the west side of the operation. LB sherds predominate as we remove sods. Upon completion of the sod removal we began an additional 10cm clearing on the eastern half of the operation. A new locus was identified on the western edge (E501) at a depth of 20cm below boulder (immovable stone). Clearing will continue tomorrow concurrent with the following of this new level.